# MMSearch-Plus: A Simple Yet Challenging Benchmark for Multimodal Browsing Agents

### v1

@@ Abstract
Large multimodal language models (MLLMs) are increasingly deployed as web agents, yet many multimodal browsing benchmarks can be solved by shallow, fixed workflows that lean on high-recall image search and nearby text—masking the genuinely multimodal challenges of fine-grained visual reasoning, provenance verification, and long-horizon tool use. We introduce MMSearch-Plus, a benchmark of 311 tasks that highly demands multimodal understanding while preserving the difficulty profile of strong text-only browsing suites. Each item is constructed to contain multiple weak, localized visual signals that must be extracted, propagated through iterative text–image search, and cross-validated under retrieval noise before answering. Our curation procedure, Spatial–Temporal Extrapolation, seeds questions whose answers require extrapolating from spatial cues (micro-text, part-level appearance, layouts, signage) and temporal traces (broadcast overlays, seasonal context) to out-of-image facts such as events, dates, and venues. We provide a model-agnostic agent framework with browsing tools, and evaluate a range of closed and open MLLMs. The strongest agent (o3) attains 15.1% without search and 36.0% accuracy with rollout under our framework, while a strong open-source model (Qwen-2.5-VL-72B-Instruct) achieves 0.0% without search and 6.9% after 20 rounds of search. Beyond answer accuracy, we assess bounding-box production and cropped-image search, and conduct an error analysis that surfaces failures in source verification, part-based reasoning, and long-horizon planning.
Project Page: https://mmsearch-plus.github.io
@@ 1 Introduction
Building on these insights, we argue that a strong multimodal benchmark should: (i) require fine-grained, exhaustive visual reasoning that compels models to mine subtle, localized cues rather than rely on a single salient (“dominant”) entity that occupies most of the image; and (ii) enforce provenance and source verification under retrieval noise—i.e., discriminating authentic sources when image results are noisy or conflicting and validating images embedded in webpages. Concretely, the benchmark should stress robust cross-modal evidence gathering, careful verification, and resilience to near-duplicates. For example, two football matches from adjacent seasons may share nearly identical visuals (kits, stadium angles). Correctly answering requires inspecting localized markers (e.g., patch details, sponsor variants, scoreboard typography) and cross-checking dates and venues against trustworthy sources. In real applications, difficult multimodal information-seeking tasks exhibit three recurring challenges that our benchmark targets:
@@
    Professional multimodal handling and long, tool-augmented chains. Agents must extract answers from heterogeneous media (images, documents, videos) across the web, sustain long reasoning/tool-use chains, and invoke visual tools programmatically (e.g., cropping, OCR) when helpful, akin to advanced agents such as o3/GPT-5-Thinking.
Our benchmark.
@@
1.1 Contributions
Our contributions are as follows:
@@ 1.1 Contributions
  • We introduce the curation procedure Spatial-Temporal Extrapolation to construct difficult multimodal search questions that require extrapolating from localized spatial cues and temporal signals to out-of-image facts (e.g., events, dates, venues), thereby stressing provenance verification under retrieval noise.
  • We develop a custom search agent framework and provide a comprehensive evaluation of current MLLMs in this setting, revealing gaps and opportunities for agents capable of sophisticated, tool-augmented multimodal reasoning.
A concurrent work, BrowseComp-VL (Geng et al., 2025), increases difficulty mainly by expanding the text search space in a BrowseComp-like manner, while the image often reduces to a single identifiable entity that is straightforwardly searchable and quickly disposable. More concretely, BrowseComp-VL from (Geng et al., 2025) is built by first creating multi-hop text QA (BrowseComp-style, with entity obfuscation) and then converting to VQA by replacing the explicit entity mention with an image of that entity (via web image retrieval). As a result, many problems primarily hinge on text search and webpage visiting after an initial visual anchoring step, rather than sustained fine-grained visual reasoning. Another concurrent work,  (Li et al., 2025b), is related. Our contributions differ in (a) data sources; (b) a general search framework that supports any MLLM; and (c) an analysis of whether thinking with images/cropping helps current MLLMs excel on our benchmark.
@@ 2.1 Spatial-Temporal Extrapolation
A central challenge in BrowseComp-like benchmarks arises from the large intermediate search space induced by soft, fuzzy constraints. This requires agents to perform non-trivial cross-validation and identify the correct target. In designing our benchmark, rather than remixing existing text-only datasets, we aim to construct problems that naturally expand the search space during multimodal information seeking, thereby testing an agent’s ability for strategic planning and uncertainty-aware reasoning in dynamic environments.
Inspired by Geoguessr-style tasks, our problems are anchored on real-world events. Agents must piece together fragmented visual information to identify the underlying source event. Task difficulty is modulated by varying the richness of both visual clues and textual context. Even a single visual fragment can expand the search space dramatically, requiring careful comparison with retrieved content and cross-validation against other multimodal evidence. In more difficult cases, this mirrors human cognition: the agent must iteratively generate hypotheses, verify them against internal knowledge or retrieved content, and refine its reasoning chain across interleaved text and images. Such processes result in extended trajectories that demand robust contextual understanding.
@@
Once an event is identified, we formulate questions that probe its metadata or chain together multi-hop queries. To further elevate difficulty, we introduce Spatial-Temporal Extrapolation. Instead of asking what is directly visible, we query what is contextually implied but physically absent, compelling reasoning beyond the pixels to reconstruct the broader event. Spatial extrapolation targets unseen entities—individuals off-screen, facing away, or partially obscured—while temporal extrapolation probes events preceding or following the depicted moment. This design forces agents to first localize the event precisely (e.g., time, match, or episode), and then retrieve and reason over wider contextual knowledge from diverse sources.
@@ 2.2 Data Collection
Building on the Spatial-Temporal Extrapolation framework, we instantiated the data collection process as follows.
We began with a set of predefined categories and generated search keywords using either LLM assistance or manual curation. Data were primarily collected from three sources: YouTube, Bilibili, and ArXiv. Keywords from categories such as Vlog were searched on YouTube and Bilibili, while those in the Academic category were searched on ArXiv, where relevant research papers (PDFs) were retrieved. We treat both videos and papers as events.
For videos, human annotators manually selected keyframes suitable for constructing questions. The keyframes underwent deanoymouization for sensitive content. For papers, annotators extracted figures or tables, capturing them via screenshots. A screenshot was considered suitable if it (i) contained vague or noisy information that challenged perception, or (ii) depicted less familiar entities requiring external retrieval to answer associated questions. This ensured that events were either (a) novel, i.e., published after model knowledge cutoffs, or (b) rarely represented in training corpora, consistent with the principles in (Jiang et al., 2024).
@@
  • Image perturbation: Critical visual tokens (e.g., unique logos, labels) were blurred or masked.
  • Text sanitization: Explicit metadata (e.g., titles, author names) was redacted while maintaining semantic content.
@@ 2.3 Dataset Statistics
Metric Value
#Total Entries 311
#Total Entries w/ Images 311
Avg #Images per Entry 1.418
Avg #Words in Question 12.273
Avg #Words in Answer 3.840
#Primary Categories 8
#Secondary Categories 43
Figure 3: (a) Distribution of primary categories in MMSearch-Plus. (b) Summary of statistics.
Figure 3 depicts the distribution of MMSearch-Plus tasks across eight primary categories, which is approximately balanced. Each task includes at least one image, with additional images provided when a single view is insufficient. The benchmark primarily comprises short-answer questions, with answers averaging 3.8 words. In hindsight, we labeled the difficulty level of a task as Easy if it can be answered correctly without search or with only image search by o3 or Gemini. Otherwise, the task is labeled as Difficult (interchangeably Hard).
@@ 2.4 Benchmark Metric and Evaluation
We adopt accuracy as the primary evaluation metric. For each question, a set of acceptable answers is provided. We use LLM-as-a-judge to compare model predictions against this set, deeming an answer correct if it matches any acceptable variant. Validation through manual inspection showed complete agreement between human judgment and LLM assessment.
@@ 2.5 Search Agent Framework
We leveraged SerpApi to support both text and image search. The framework is compatible with sub-image cropping and sub-image search for fine-grained analysis. However, due to the limited grounding ability of the tested MLLMs, we did not incorporate sub-image cropping in the main experiments.
The framework incorporated several optimizations:
  • Search result caching: For each image in MMSearch-Plus, we cache the corresponding image-search results (i.e., the ranked URLs and thumbnails returned by the engine). This reduces latency and token cost during repeated agent steps and evaluation, and it does not deterioriate reproducibility because the top image-search results are relatively stable across time in our setting.
  • Top-$K$ search results retrieval: The most relevant search results determined by the search engine are passed to the MLLMs.
  • Content extraction: Retrieved webpages are summarized by an MLLM into web_info and related_info, discarding boilerplate HTML to reduce token cost. web_info contains information relevant to the question, whereas related_info captures the relevance between a webpage’s thumbnail and the query image.
  • Contextualized evaluation: The agent is always provided with the full conversation history and prior reasoning steps.
Table 1: End-to-end results on the MMSearch+ benchmark, across search modes. All numbers are accuracy (%).
@@
Closed-source LMMs
@@
Open-source LMMs
Qwen-2.5-VL-72B-Instruct
@@ 3.1 Experiment Setup
We evaluate both closed-source and open-source MLLMs on our benchmark. Closed-source models include Gemini-2.5-Pro (hereafter Gemini), o3, and GPT-5, representing the current frontier of proprietary multimodal systems. On the open-source side, we focus on Qwen-2.5-VL-72B-Instruct (hereafter Qwen), a state-of-the-art vision–language model trained at large scale. All models are used in their publicly released inference versions without any task-specific fine-tuning.
@@ 4 Results
Baseline Performance (Without Search)
Without any external retrieval, the open-source model (Qwen) achieved 0.0% across all categories, indicating that the benchmark is intentionally adversarial to parametric-only knowledge. Closed-source baselines fare better but are still far from saturation: GPT-5 at 10.3%, Gemini-2.5-Pro at 10.6%, and o3 at 15.1%. The category breakdown reveals why: even for the strongest no-search model (o3), questions that likely require long-tail facts or fine-grained entity disambiguation remain difficult (e.g., Games 3.2, Vlog 5.3, Acad. 6.0), whereas more visually grounded or canonical domains such as Geo. (31.2) and the Easy split (50.0) are comparatively tractable. This spread suggests that, on MMSearch+, in-context parametric knowledge alone is insufficient: many items appear to require explicit grounding to web evidence, especially where named entities, screen text, or episodic, pop-culture–style facts dominate.
Figure 4: Accuracy of o3 and Gemini-2.5-Pro across 3 search modes.
Effect of Image Search
A single round of image search provides a consistent, modest lift: on average +4.2 points for o3 (from 15.1 to 19.3), +6.1 for GPT-5 (10.3 to 16.4), and +5.8 for Gemini-2.5-Pro (10.6 to 16.4). Qwen also benefits, jumping from 0.0 to 13.5 (+13.5). Overall, the pattern implies that a single, reasonably relevant image can unlock coarse disambiguation for Qwen but leaves complex, compositional cases unresolved.
Full Rollout
Allowing iterative search yields a step-change in performance. o3 reaches 36.0% (absolute +20.9 over no-search), and Gemini-2.5-Pro reaches 23.8% (+13.2). The pattern aligns with tasks that require progressive narrowing: o3’s largest category deltas come in Acad. (+44.0; 6.0 $\rightarrow$ 50.0), Vlog (+36.8), and Tech (+30.5), where multi-hop browsing can surface specific frames, quotes, or specs. Gemini also shows substantial lifts in Geo. and Music (both +23.5), Tech (+19.4), and Film/TV (+15.0). By contrast, Qwen regresses under our rollout controller: its average drops from 13.5 (image search) to 6.1 (–7.4). This suggests that iterative search is not plug-and-play for Qwen: without model-specific tuning of the rollout policy (e.g., hop budgeting, snippet ranking, stopping criteria) and/or stronger web-reading robustness, extra hops can introduce distractors faster than they add evidence.
@@ 5 Analysis
5.1 Rollout Trajectory Statistics
Figure 5 provides an overview of reasoning-trajectory statistics across three MLLMs (o3, Gemini, Qwen): the left column reports the distribution of text- and image-search calls per conversation, while the right column relates assistant word count to the number of search calls, stratified by correctness. All statistics are collected in full rollout mode. From the figure, we summarize:
  • More calls when wrong. Incorrect trajectories typically include more text-search calls than correct ones, indicating that uncertainty triggers extra querying without surfacing decisive evidence.
  • Gemini under-explores. Despite a budget of $20$ calls, Gemini usually issues fewer than $5$ search calls per trajectory; text-search medians hover near 011 and image-search is near zero for both correct and incorrect cases. When it does search, response length grows mildly with call count, suggesting reliance on parametric knowledge rather than iterative retrieval.
  • Little link between calls and verbosity. Across models there is no robust monotonic relationship between the number of calls and the length of the assistant’s response. Trends are model-dependent: o3 shows a slight negative slope, Gemini a mild positive slope, and Qwen’s incorrect trajectories become more verbose with more calls while correct ones do not.

### v2

# MMSearch-Plus: Benchmarking Provenance-Aware Search for Multimodal Browsing Agents

@@ Abstract
Existing multimodal browsing benchmarks often fail to require genuine multimodal reasoning, as many tasks can be solved with text-only heuristics without vision-in-the-loop verification. We introduce MMSearch-Plus, a 311-task benchmark that enforces multimodal understanding by requiring extraction and propagation of fine-grained visual cues through iterative image–text retrieval and cross-validation under retrieval noise. Our curation procedure seeds questions whose answers require extrapolating from spatial cues and temporal traces to out-of-image facts such as events, dates, and venues. Beyond the dataset, we provide a model-agnostic agent framework with standard browsing tools and a set-of-mark (SoM) module, which lets the agent place marks, crop subregions, and launch targeted image/text searches. SoM enables provenance-aware zoom-and-retrieve and improves robustness in multi-step reasoning. We evaluated closed- and open-source MLLMs in this framework. The strongest system achieves an end-to-end accuracy of 36.0%, and integrating SoM produces consistent gains in multiple settings, with improvements up to +3.9 points. From failure analysis, we observe recurring errors in locating relevant webpages and distinguishing between visually similar events. These results underscore the challenges of real-world multimodal search and establish MMSearch-Plus as a rigorous benchmark for advancing agentic MLLMs.
Project page: https://mmsearch-plus.github.io/
@@ 1 Introduction
Figure 1: Three multimodal reasoning paradigms: 1. Without search: an MLLM reasons with its internal knowledge to answer a factual visual question (VQA); 2. Whole image search only: an MLLM combines its internal knowledge and the provided search results of an VQA’s image; 3. MMSearch-Plus agentic framework: an MLLM can call a set of visual and search tools freely to extracts fine-grained visual clues and search with precision. A real example from MMSearch-Plus is provided in Figure 8.
@@
Building on these insights, we argue that a strong multimodal benchmark should: (i) require fine-grained, exhaustive visual reasoning that compels models to mine subtle, localized cues rather than rely on a single salient (“dominant”) entity that occupies most of the image; and (ii) enforce source verification under retrieval noise—i.e., discriminating authentic sources when image results are noisy or conflicting and validating images embedded in webpages. Concretely, the benchmark should stress robust cross-modal evidence gathering, careful verification, and resilience to near-duplicates. For example, two football matches from adjacent seasons may share nearly identical visuals (kits, stadium angles). Correctly answering requires inspecting localized markers (e.g., patch details, sponsor variants, scoreboard typography) and cross-checking dates and venues against trustworthy sources. In real applications, difficult multimodal information-seeking tasks exhibit three recurring challenges that our benchmark targets:
@@
    Professional multimodal handling and long, tool-augmented chains. Agents must extract answers from heterogeneous media (images, documents, videos) across the web, sustain long reasoning/tool-use chains, and invoke visual tools programmatically (e.g., cropping, OCR) when helpful, akin to advanced agents such as o3.
@@
Our key contributions can be summarized as follows:
@@
  • We introduce the curation procedure Spatial-Temporal Extrapolation ($\xi$ 2.1) to construct difficult multimodal search questions that require extrapolating from localized spatial cues and temporal signals to out-of-image facts.
  • We develop a general, model-agnostic search agent framework that interleaves text/image search with a Set-of-Mark zoom-in pathway ($\xi$ 2.5). SoM supplies human-verified marks and region-seeded retrieval (zoom_in/image_search), enabling sustained fine-grained reasoning; we ablate SoM to quantify how cropping and thinking with images affect current MLLMs.
A concurrent work, BrowseComp-VL (Geng et al., 2025), increases difficulty mainly by expanding the text search space in a BrowseComp-like manner, while the image often reduces to a single identifiable entity that is straightforwardly searchable and quickly disposable. Another concurrent work, MM-BrowseComp (Li et al., 2025b), is related. Our contributions differ in (a) data sources and (b) a general search framework that also enables thinking with images via Set-of-Mark.
@@ 2.1 Spatial-Temporal Extrapolation
Summary. We test an agent’s ability to infer facts that are not literally present in the pixels or the instant shown, requiring it to recover the broader event by combining partial visual cues with retrieval and reasoning across time and space. This emphasis discourages purely perceptual shortcutting and directly measures uncertainty-aware planning, precise event localization, and evidence-backed browsing.
A central challenge in BrowseComp-like benchmarks is the ballooning intermediate search space induced by soft, fuzzy constraints, which demands non-trivial cross-validation to identify the correct target. Rather than remixing text-only corpora, we construct multimodal problems that naturally expand this search space: from a handful of visual fragments tied to real-world events (in the spirit of GeoGuessr AB (2013)), the agent must hypothesize the underlying source event and verify it against retrieved evidence. Difficulty is modulated by the richness of visual and textual cues; even a single crop can widen the candidate set dramatically, eliciting iterative hypothesis–test–refine behavior and long trajectories that require robust context tracking.
@@
Once the source event is pinned down, our questions probe metadata and support multi-hop reasoning. To elevate difficulty, we introduce Spatial-Temporal Extrapolation: instead of querying what is directly visible, we ask what is contextually implied but physically absent, compelling reasoning beyond the frame and moment to reconstruct the broader situation. Spatial extrapolation targets unseen entities (e.g., off-screen participants, individuals facing away, occluded signage), while temporal extrapolation probes events immediately before or after the depicted moment (e.g., lineups, the next play/goal, or the subsequent segment/episode). Solving these requires agents to (i) precisely localize the event (time, match, or episode), and (ii) retrieve and integrate wider contextual knowledge from diverse sources to produce an evidence-grounded answer.
@@ 2.2 Data Collection
Building on the Spatial–Temporal Extrapolation framework, we instantiated the data collection process as follows.
We began with a set of predefined categories and generated search keywords using a mix of LLM assistance and manual curation. Data were sourced from publicly accessible video-sharing platforms and open-access scholarly repositories (i.e., arXiv). Queries for consumer video content (e.g., Vlog) were issued to video platforms, while those in the Acad. category were issued to arXiv to retrieve relevant research papers (PDFs). All materials are used only for research. We also mask sensitive content (e.g., faces and vehicle license plates). In our pipeline, both videos and papers are treated as events.
For videos, human annotators manually selected keyframes suitable for constructing questions. The keyframes underwent anonymization for sensitive content. For papers, annotators extracted figures or tables, capturing them via screenshots. A screenshot was considered suitable if it (i) contained vague or noisy information that challenged perception, (ii) depicted less familiar entities requiring external retrieval to answer associated questions, and (iii) could not be searched with Google image search directly. This ensures that most, if not all, screenshots are not in the training dataset, and the essential information for answering the question is neither in the training corpora nor directly retrievable via a direct image search.
@@
  • Image perturbation: Critical visual subregions were blurred or masked.
@@ 2.3 Dataset Statistics
Statistic Number
Total questions w/ images 311
Primary categories 8
Secondary categories 43
Unique images 441
Unique questions 279
Unique answers 306
Maximum question length 42
Maximum answer length 20
Average question length 12.5
Average answer length 3.7
Difficulty distribution
- Easy 94 (30.2%)
- Difficult 217 (69.8%)
Figure 3: (a) Distribution of primary categories in MMSearch-Plus, with light and dark inner sectors representing the easy and difficult splits within each category. (b) Summary of statistics.
Figure 3 depicts the distribution of MMSearch-Plus tasks across eight primary categories, which is approximately balanced. Each task includes at least one image, with additional images provided when a single view is insufficient. The benchmark primarily comprises short-answer questions, with answers averaging 3.7 words. In hindsight, we labeled the difficulty level of a task as Easy if it can be answered correctly without search or with only image search by o3 or Gemini-2.5-Pro. Otherwise, the task is labeled as Difficult (interchangeably Hard).
@@ 2.4 Benchmark Metric and Evaluation
We adopt accuracy as the primary evaluation metric. For each question, a set of acceptable answers is provided. We use LLM-as-a-judge to compare model predictions against this set, deeming an answer correct if it matches any acceptable variant. Validation through manual inspection showed complete agreement between human judgment and the assessment of GPT-4o.
@@ 2.5 Search Agent Framework
Overview.
We implement a model-agnostic web agent that interleaves text search, image search, and a zoom-in tool. Ranked results are obtained via SerpApi ; the surrounding framework (i.e., the system that includes the MLLM agent) performs lightweight post-processing to summarize pages, relate images to the query, and manage multi-round context. To reduce latency and token cost across repeated steps/evaluations, the framework caches image-search returns (ranked URLs/thumbnails) and their MLLM summaries per benchmark image. The framework maintains a threaded conversation state (tool calls, cropped views, summaries, and hypotheses) to support long-horizon, contextualized reasoning. It is compatible with sub-image cropping and sub-image search for fine-grained analysis.
Refer to caption Q: Who assisted this goal? A: Philipp Lahm
Figure 4: Example of an SoM-overlaid task image.
Text and image search.
At any step, the agent may invoke text or image search. SerpAPI returns ranked candidates, from which the framework forwards a small set of top results to the model (images are summarized and cached; webpages are summarized). Retrieved webpages are summarized by an MLLM (Gemini in our runs) into two fields:
  • web_info: task-conditioned semantic summaries of page content that may support answering the current question;
  • related_info: evidence linking a result’s thumbnail/lead image(s) to the query image (e.g., matching signage, layouts, or micro-text).
These summaries are designed to compress interaction history for use within limited MLLM context windows; metadata such as URL/title is readily handled by SerpAPI. All retrieved and generated artifacts are integrated into the conversation state to enable contextualized evaluation across turns.
Zoom-in via Set-of-Mark (SoM).
To enable precise, provenance-aware inspection, we augment the agent with a Set-of-Mark (SoM) module (Yang et al., 2023). In this work, we refer to SoM as a list of human-verified bounding boxes per task image (examples in Fig. 9). In the first turn, we present both the raw image and an overlaid version with box outlines and indices, avoiding occlusion while allowing cross-reference.
Table 1: End-to-end results on the MMSearch+ benchmark, across search modes. All numbers are accuracy (%). Darker blue means higher accuracy. The best and second-best methods under each setting are shown in bold and underline, respectively.
@@
Human
Browser 22.8 20.3 25.9 20.0 25.0 19.4 16.1 31.6 35.3 34.0 18.0
Closed-source
@@
Full Rollout + SoM 27.7 40.6 22.2 24.0 25.0 33.3 19.4 15.8 29.4 54.3 16.1
Open-source
Qwen-2.5-VL-72B
@@ 3 Experiment
We evaluate both closed-source and open-source MLLMs on our benchmark. Closed-source models include Gemini-2.5-Pro (hereafter Gemini(Team & DeepMind, 2025), o3 (OpenAI, 2025b) and GPT-5 (OpenAI, 2025a), representing the current frontier of proprietary multimodal systems. On the open-source side, we focus on Qwen-2.5-VL-72B-Instruct (hereafter Qwen(Bai et al., 2025). Expert human annotators were instructed to work with a Chrome browser for at least 10 minutes for each task, and capped at 20 minutes.
@@
  4. (d)
    Full Rollout + SoM. Images overlayed with Set-of-Mark are provided to MLLMs. In addition to capabilities provided in Full Rollout, models can (a) call the zoom-in tool to examine a subregion of an image, (b) call the image search tool to search with the subregion. The maximum number of tool calls is 20.
@@ 4 Results
Under No Search, open-source Qwen scored 0.0%, confirming that MMSearch+ resisted parametric shortcuts. Closed-source models remained far from saturation (GPT-5 10.3, Gemini-2.5-Pro 10.6, o3 15.1). With Image Search (o3 +4.2, GPT-5 +6.1, Gemini +5.8; Qwen +13.5), models gained from coarse disambiguation, yet multi-hop composition remained the main blocker. With iterative image and text search, Full Rollout produced large gains: o3 36.0 (+20.9 over no-search) and Gemini 23.8 (+13.2), while Qwen regressed to 6.1 (-7.4 over image-search). This indicates that the external information retrieved from search tools is essential to answering the questions correctly. Notably, easy split accuracy did not improve from Image Search to Full Rollout. We attribute this to two factors:
  1. 1.
    Label vs. retrieval difficulty. In MMSearch+, “easy” meant higher likelihood of being in-model or solvable by a direct image hop, not that multi-step retrieval was trivial.
  2. 2.
    Policy overhead and distractors. Full rollout added branching (tool choice, query rewriting). For easy items solvable by a single image search, extra hops often injected distractors (off-topic pages, SEO noise), increasing drift from the correct evidence.
Overall, rollout helped where progressive narrowing was necessary (academic, tech, media queries), but on easy items extra freedom reduced precision.
Adding SoM (Full Rollout + SoM) gave consistent, targeted gains: o3 37.7 (+1.7), Gemini 27.7 (+3.9), Qwen +1.0 over Full Rollout. Set-of-Mark anchored reasoning on fine-grained visual cues (micro-text, logos, key entities) before expanding the search space.
@@ 5 Analysis
We present a list of observations (Section 5.1) and an error analysis (Section 5.2).